OpenStudy (anonymous):

If f(x) = −4x − 1, find f(−2).

hartnn (hartnn):

just put x=-2 -4(-2)-1 = ??

OpenStudy (anonymous):

i got -7

hartnn (hartnn):

nopes, -4(-2)=8 and 8-1 = 7 not -7